# “Feeling that I was Collaborating with Them:” A 20-year Systematic Literature Review of Social Virtual Reality Leveraging Collaboration

### v2

@@ 4.2.1. Individual Perception and Experience of Collaboration in Social VR
“Am I Really There?” Sense of Presence as a Foundation for Immersive Collaboration.
This sub-theme focuses on individual users’ sense of presence—the subjective feeling of “being there” within a 3D virtual environment (Triberti et al., 2025)—as a foundational condition for collaboration in Social VR. Distinct from social presence, which refers to the perception of others in the environment, sense of presence emphasizes the users’ personal immersion in the virtual space. This experience is often shaped by how intuitively users’ actions and intentions are translated into the virtual environment (Oh et al., 2018; Olaosebikan et al., 2022).
Several studies in our corpus indicated that Social VR can enhance a strong sense of presence by allowing users to act and interact in more natural and embodied ways (Olaosebikan et al., 2022; Khojasteh and Won, 2021; Mei et al., 2021). For instance, Olaosebikan et al. (Olaosebikan et al., 2022) found that scientists collaborating in VR reported a greater sense of immersion and presence compared to those using traditional video conferencing platforms. Participants described their VR experiences as attending a live presentation, highlighting the engaging and spatial qualities of immersive environments (Olaosebikan et al., 2022).
Similarly, (Khojasteh and Won, 2021) conducted a five-session longitudinal study in which users repeatedly collaborated in VR with different partners. Over time, participants reported increased feelings of presence, suggesting that growing familiarity with the virtual environment, task, and interaction patterns contributes to a more immersive experience (Khojasteh and Won, 2021).
Mei et al. (Mei et al., 2021) explored how a co-design activity in VR supported ideation and task engagement. Participants reported that being able to spatially manipulate and visualize design elements within a shared 3D virtual environment enhanced their sense of “being in the task.” Their findings highlight how interactive design tools can strengthen users’ presence through embodied interaction.
In sum, these studies suggest that fostering a strong sense of presence is necessary for effective collaboration in Social VR. When the applications support natural movement, spatial awareness, and meaningful shared activities, users are more likely to feel grounded in the virtual environment and interact intuitively with others. Over time, repeated exposure and growing familiarity with the virtual space can deepen users’ immersion, enabling collaborators to engage more fluidly and confidently.
@@
This sub-theme examines how avatar design influences users’ experiences in Social VR. We found that multiple articles focused on both how users perceive their own virtual bodies (i.e., self-perception) and how they believe others perceive them. Their main findings show that avatar embodiment plays a key role in how users express themselves, interpret social cues, and participate in collaborative settings. The visual and behavioral fidelity of avatars impacts users’ sense of identity and presence, influencing their confidence, expressiveness, and overall engagement during collaboration.
@@
The articles in this sub-theme demonstrated that avatar characteristics— such as gender, realism, level of abstraction (i.e., how simple or complex the avatar appears), and humanoid forms—significantly shape users’ experiences (Osborne et al., 2023; Torro and Pirkkalainen, 2023; Yang et al., 2024). Gender representation emerged as a recurrent aspect across these studies. For instance, Yassien et al. (Yassien et al., 2021) found that having all the group members’ avatars with the same gender (e.g., all-male or all-female) improved collaboration quality and mutual support, even when users’ real-world gender was different. Similarly, Yang et al. (Yang et al., 2024) demonstrated that users represented by non-human avatars (e.g., fox, robot, duck) reported greater comfort and equal participation than when using communication formats where their real-world appearance was visible, such as face-to-face or video conferencing settings, suggesting that abstract representations may alleviate social pressure and enable more balanced collaboration.
@@
Other studies emphasized the importance of avatar customization and stylization when collaborating in Social VR (Osborne et al., 2023; Freeman et al., 2020). Osborne et al. (Osborne et al., 2023) showed that allowing users to personalize their avatars’ appearance enhanced engagement with the rest of the group. Similarly, Freeman et al. (Freeman et al., 2020), in a descriptive study involving 30 semi-structured interviews (Freeman et al., 2020), demonstrated that enabling self-representation in Social VR supports identity exploration, particularly for users navigating gender or sexual identity. Participants often described their avatars as extensions of themselves, using them to explore self-image in ways that felt both safe and empowering (Freeman et al., 2020).
@@
However, some articles also showed that increased realism in avatar design can complicate users’ sense of self-identification by creating a mismatch between the avatar’s appearance and the user’s internal self-image. Two studies addressed how the “uncanny valley” effect—which occurs when avatars look nearly human but still exhibit subtle imperfections in appearance or movement—made participants feel unnatural or unsettled during the collaboration experience (Osborne et al., 2023; Torro and Pirkkalainen, 2023). Osborne et al. (Osborne et al., 2023) and Torro et al. (Torro and Pirkkalainen, 2023) reported that these unsettling avatars disrupted users’ immersion and hindered their natural interactions. While humanoid avatars can promote familiarity and trust, these findings underscore the importance of addressing the uncanny valley when designing avatars for collaborative experiences in Social VR.
@@
Several articles examined how the fidelity of avatars’ motion and expression influences users’ perceptions of being accurately represented in their social interactions. Kimmel et al. (Kimmel et al., 2024) and Yassien et al. (Yassien et al., 2020) found that full-body avatars with realistic tracking and expressive features (e.g., gaze direction, facial movement) helped users feel more accurately represented. Similarly, Kocur et al. (Kocur et al., 2020) further examined how users believed they were perceived by their teammates in VR. They found that mismatches between how individuals wanted to present themselves and how they believed others perceived them—often due to avatar limitations—undermined users’ confidence and reduced their engagement in group collaboration.
@@
In summary, these articles demonstrate that avatars are not merely cosmetic features but critical components of users’ collaborative experiences in VR. They influence how individuals perceive themselves, their social roles, expressions, and feelings of being acknowledged within shared spaces. As the articles in this sub-theme illustrate, design choices regarding avatar realism, abstraction, gender, and expressiveness can deeply affect users’ comfort, sense of inclusion, and confidence during collaborative tasks in VR.
@@ 4.2.2. Team Dynamics and Collaboration in Virtual Environments
This sub-theme explores how social presence—the sense of “being with others” in a shared virtual environment (Oh et al., 2018)—facilitates effective team collaboration in Social VR. While the sense of presence refers to a user’s perception of immersion, social presence refers to the awareness of and connection to teammates. It addresses the feeling that others are truly co-located and responsive within the virtual space.
Several articles emphasized that social presence is a foundational social-cognitive state that enables key collaborative behaviors such as turn-taking, mutual coordination, and shared understanding (Olaosebikan et al., 2022; Dey et al., 2024). In a systematic review of Social VR teamwork studies, Dey et al. (Dey et al., 2024) found that social presence strongly influences team performance, citing increased mutual awareness, better turn coordination, and stronger engagement as key team outcomes.
@@
A key component of social presence in VR is mutual attention, which is the ability to perceive others’ attention and actions in real-time (Darbar et al., 2024; Bovo et al., 2022; Ide et al., 2020). For example, Darbar et al. (Darbar et al., 2024) introduced ‘GazeMolVR,’ a system for molecular visualization in collaborative VR that incorporated shared eye-gaze cues. Their findings showed that real-time gaze sharing improved the feeling of connectedness and eased conversational flow, particularly in distributed teams. When collaborators could see where others were looking, they adjusted their own attention more fluidly, making the interaction feel more natural and co-present (Darbar et al., 2024).
@@
The articles also show that social presence is strongly connected to nonverbal awareness and emotional connection. When avatars or systems convey cues such as gaze direction, pointing, or simple gestures, teammates can more easily interpret each other’s intentions (Bovo et al., 2022; Ide et al., 2020). For example, studies such as Bovo et al. (Bovo et al., 2022) and Ide et al. (Ide et al., 2020) implemented visual aids, including cones of vision and symbolic gestures, to help users perceive their collaborators’ focus and actions. These features not only added expressiveness but also enhanced social presence by increasing mutual visibility and acknowledgment, helping users feel noticed and understood by others.
@@
One article highlighted that social presence enhances coordination during multi-party interactions (Adkins et al., 2024). In their evaluation of audio and interaction modalities, Adkins et al. (Adkins et al., 2024) found that teams using spatial audio, compared to non-spatial conditions, experienced smoother turn-taking and reported feeling more “in sync” with one another. These findings suggest that social presence is shaped not only by visual or avatar-based cues but also by auditory and behavioral alignment that supports team coordination.
@@
One last article noted that weak or underdeveloped social presence negatively affects collaboration. In their evaluation of a co-design task, Mei et al. (Mei et al., 2021) found that the absence of avatar reactivity and expressive behavior made interactions feel more like a phone call rather than a shared, embodied collaboration. This diminished sense of social presence reduced user engagement and made communication feel less natural, highlighting the importance of social presence for enabling teamwork.
@@
In summary, these articles demonstrate that social presence serves as the social glue that makes collaboration in Social VR feel natural and connected. The articles show how it fosters mutual attention, real-time responsiveness, emotional connection, and spatial alignment, all of which contribute to a strong sense of togetherness. Ultimately, social presence enables team members to perceive and respond to one another’s actions in real time, supporting cohesive and authentically shared collaboration in VR.
@@
This sub-theme includes articles that examine how users evaluate the authenticity, seriousness, and naturalness of their collaborative experiences in Social VR. The focus is on whether immersive VR environments support interactions that are not only technically possible but also feel socially valid, natural, and meaningful. Beyond the technical functionality, these articles highlight that users carry expectations shaped by physical-world norms, emotional cues, and social dynamics, which influence how “real” collaboration feels in Social VR (Torro and Pirkkalainen, 2023; Maloney and Freeman, 2020; Osborne et al., 2023).
@@
Building on this, Maloney et al. (Maloney and Freeman, 2020) interviewed long-term Social VR users to understand what made virtual interactions feel meaningful. Participants reported that their collaborations felt more authentic when gestures, feedback, and shared activities aligned with everyday social norms. Many of them expressed frustration with unrealistic avatar features—such as exaggerated cartoon-like hands or faces that could not show realistic emotions—which made the interactions feel artificial and less socially authentic. What users desired was not just visual realism, but interaction fidelity, which translated into responsiveness and expressiveness that matched the tone and rhythm of real-world collaboration.
Other articles in this sub-theme also indicated that misaligned avatar behavior can disrupt the perceived authenticity of collaboration. For example, Osborne et al. (Osborne et al., 2023) noted that technical glitches, such as avatar duplication or unnatural expressions, made interactions feel artificial and awkward. In contrast, when avatars moved and responded in socially expected ways, participants perceived the collaboration as smooth and believable, even when the avatars were abstract in design (Osborne et al., 2023).
@@
One study examined how to reinforce the role of realism in a Social VR workplace by eliminating fantastical elements. Park et al. (Park et al., 2024) removed the teleportation feature and required users to navigate through halls and elevators, mirroring how people would walk in a physical office. While this approach increased social encounters and enhanced users’ social presence, it also introduced usability challenges, revealing a tension between realism and practicality (Park et al., 2024). The findings of this article echo broader concerns about how much realism is necessary to foster authentic collaboration without compromising the unique affordances of VR.
Lastly, one article (Dey et al., 2024) showed that Social VR who first engaged in an emotionally intense solo experience approached a subsequent group task with more seriousness and coordination. The realism of the initial simulation made the following collaboration feel more consequential and authentic. Dey et al.’s findings (Dey et al., 2024) suggest that initial immersive experiences can reshape how users’ perceptions of group work in Social VR, making virtual collaboration feel more meaningful, important, and socially grounded.
In summary, these articles demonstrate that the authenticity of collaboration in Social VR is not just a result of visual fidelity or technical features, but on how well interactions mirror the rhythms, responsiveness, and emotional tone of real-world teamwork. When Social VR users feel that their actions have an impact and that their collaborators are genuinely engaged, the experience transcends the simulation and becomes a psychologically meaningful, socially credible form of collaboration.
@@
This sub-theme explores how trust, an essential yet frequently overlooked foundation of effective collaboration, is developed and sustained in Social VR. In co-located settings, trust often emerges through informal interactions and subtle nonverbal cues. However, these trust-building mechanisms are often limited or absent in remote communication settings. Nevertheless, the reviewed articles show that Social VR can foster interpersonal trust through its distinctive affordances, including embodied interaction, avatar behavior, and shared spatial context (Torro and Pirkkalainen, 2023; Yassien et al., 2020).
Torro et al. (Torro and Pirkkalainen, 2023) developed a design theory explaining how Social VR can support trust-building in virtual teams. Grounded in social exchange theory, which emphasizes reciprocal interactions as the foundation of relationships (Cook et al., 2013), they proposed that trust emerges through two key pathways in Social VR: a peripheral route, driven by first impressions (e.g., avatar appearance, social cues), and a central route, based on deeper assessments of a partner’s competence, integrity, and goodwill (Torro and Pirkkalainen, 2023). The article emphasized that immersive Social VR applications designed to support both verbal and nonverbal communication can enhance trust through both routes. For example, avatars that mimic natural gestures, proximity-based conversations, and realistic movement can help promote social bonding. Moreover, features like virtual break rooms with spatial properties can support casual, unstructured interactions, which are often absent in structured video calls or text-based platforms.
Yassien et al. (Yassien et al., 2020) highlighted another key dimension of trust in Social VR: personal space. This article found that users felt safer and more connected when their virtual personal space was respected. Trust was undermined when avatars came too close or moved in ways that felt erratic or unpredictable, causing discomfort and breaking the sense of social ease. The researchers recommended designing Social VR applications with mechanisms to prevent collisions and unwanted proximity between avatars, reinforcing the importance of comfort and perceived safety in virtual interactions.
In summary, these two articles suggest that trust can be supported through both social signaling and environmental design, such as expressive avatars, respectful personal space, and opportunities for informal interaction.
@@
This sub-theme focuses on social engagement, the extent to which collaborators in Social VR remain emotionally and communicatively connected during a shared task (Xu et al., 2023). Social engagement goes beyond mere attention or performance, it reflects how present, expressive, and emotionally involved individuals feel in relation to their teammates (Xu et al., 2023). The reviewed articles identified design strategies that support behavioral cues of engagement and help sustain interpersonal connections throughout the Social VR collaboration (Le Tarnec et al., 2024; Xu et al., 2023; Moharana et al., 2023).
One strategy identified in this sub-theme involves enhancing avatar behaviors to foster social engagement. Le et al. (Le Tarnec et al., 2024) introduced an augmented avatar system that generated subtle nonverbal cues—such as head nods and posture shifts—even when users did not perform them directly. These enhancements made collaborators appear more responsive and involved, leading to more natural and engaging teamwork (Le Tarnec et al., 2024). Similarly, Xu et al. (Xu et al., 2023) studied older adults co-planning virtual travel experiences using VR and found that social engagement deepened when participants shared personal memories and life stories related to the destinations. These moments of self-disclosure fostered warmth and mutual interest, highlighting that social engagement can be supported not only through interface design but also through content that invites personal connection and emotional expression.
Another article explored physiological synchrony as a deeper measure of engagement. Moharana et al. (Moharana et al., 2023) tracked heart rate patterns during a collaborative memory game and found that teams with higher physiological synchrony reported greater feelings of connection, presence, and team flow. These findings suggest that social engagement may operate at both cognitive and emotional levels and that VR environments can help capture or even facilitate these more subtle social dynamics.
Taken together, these articles show that social engagement in VR collaboration extends beyond visual realism or task efficiency. It hinges on participants feeling seen, acknowledged, and emotionally connected, which are qualities that can be nurtured through expressive avatars, thoughtful interaction design, and moments that invite genuine interpersonal connections. Therefore, designing for sustained social engagement is essential for effective, human-centered teamwork in Social VR applications.

### v3

@@ 4.2.1. Individual Perception and Experience of Collaboration in Social VR
“Am I Really There?” Sense of Presence as a Foundation for Collaboration in Social VR
This sub-theme focuses on individual users’ sense of presence—the subjective feeling of “being there” within a 3D virtual environment (Triberti et al., 2025)—as a foundational condition for collaboration in Social VR. Distinct from social presence (i.e., perceiving others), presence describes the user’s own experience of being located in the virtual space. This experience is shaped by how intuitively users’ actions and intentions are translated into the environment (Oh et al., 2018; Olaosebikan et al., 2022; Miller et al., 2021).
Studies reported mixed effects of presence: while some found VR increased presence and task engagement compared to video calls (Olaosebikan et al., 2022; Mei et al., 2021), others observed flat presence scores despite growing familiarity (Khojasteh and Won, 2021). These inconsistencies suggest that presence depends less on time in VR and more on the availability of natural nonverbal cues. This helps explain the lack of increase: adaptation alone did not lift presence because natural nonverbal channels were missing or awkward. Consistently, Mei et al. (Mei et al., 2021) propose adding emoji-like or haptic cues to convey emotion, and report that missing facial expressions weakened users’ emotional connection. Across these papers, gaps in nonverbal expression and practical hurdles (learning curve, limited in-VR tools) made it harder to feel “there” during collaboration (Olaosebikan et al., 2022; Khojasteh and Won, 2021; Mei et al., 2021).
@@
This sub-theme examines how avatar design influences users’ experiences in Social VR. We found that multiple articles focused on both how users perceive their own virtual bodies (i.e., self-perception) and how they believe others perceive them. Their main findings indicate that avatar embodiment plays a key role in how users express themselves, interpret social cues, and engage in collaborative settings. The visual and behavioral fidelity of avatars impacts users’ sense of identity and presence, influencing their confidence, expressiveness, and overall engagement during collaboration.
@@
The articles of this corpus have demonstrated that avatar characteristics—such as gender, realism, abstraction level, and humanoid forms—significantly shape users’ experiences (Osborne et al., 2023; Shoa et al., 2023; Torro and Pirkkalainen, 2023; Yang et al., 2024; Kallioniemi et al., 2017; Li et al., 2024; Young et al., 2015). Gender representation emerged as a recurrent aspect across these studies. For instance, Yassien et al. (Yassien et al., 2021) found that having all the group members’ avatars with the same gender (e.g., all-male or all-female) improved collaboration quality and mutual support, even when users’ real-world genders differed. Similarly, Yang et al. (Yang et al., 2024) demonstrated that users represented by non-human avatars (e.g., fox, robot, duck) reported greater comfort and equal participation than when using communication formats where their real-world appearance was visible, such as face-to-face or video conferencing settings, suggesting that abstract representations may alleviate social pressure and enable more balanced collaboration.
@@
Two studies emphasized the importance of avatar customization and stylization when collaborating in Social VR (Osborne et al., 2023; Freeman et al., 2020). Osborne et al. (Osborne et al., 2023) showed that allowing users to personalize their avatars enhanced engagement with the rest of the group. Similarly, Freeman et al. (Freeman et al., 2020) demonstrated that enabling self-representation in Social VR supports identity exploration, particularly for users navigating gender or sexual identity. Participants often described their avatars as extensions of themselves, using them to explore self-image in ways that felt safe and empowering (Freeman et al., 2020).
@@
However, some articles showed that increased realism in avatar design can complicate users’ sense of self-identification by creating a mismatch between the avatar’s appearance and the user’s internal self-image. Two studies addressed how the “uncanny valley” effect—which occurs when avatars look nearly human but still exhibit subtle imperfections in appearance or movement—made participants feel unnatural or unsettled during the collaboration experience (Osborne et al., 2023; Torro and Pirkkalainen, 2023). Osborne et al. (Osborne et al., 2023) and Torro et al. (Torro and Pirkkalainen, 2023) reported that these unsettling avatars disrupted users’ sense of presence and hindered their natural interactions.
@@
Several articles examined how the fidelity of avatars’ motion and expression influences users’ perceptions of being accurately represented in their social interactions. Kimmel et al. (Kimmel et al., 2024) and Yassien et al. (Yassien et al., 2020) found that full-body avatars with realistic tracking and expressive features helped users feel more accurately represented. Similarly, Kocur et al. (Kocur et al., 2020) found that mismatches between how individuals wanted to present themselves and how they believed others perceived them—often due to avatar limitations—undermined users’ confidence and reduced their engagement in group collaboration.
@@
In summary, these articles demonstrate that avatars are not merely cosmetic features but critical components of users’ collaborative experiences in VR. They influence how individuals perceive themselves, their social roles, expressions, and feelings of being acknowledged within shared spaces.
@@ 4.2.2. Team Dynamics and Collaboration in Virtual Environments
This sub-theme explores how social presence—the sense of “being with others” in a shared virtual environment (Oh et al., 2018)—facilitates effective team collaboration in Social VR. While the sense of presence refers to a user’s perception of presence, social presence refers to the awareness of and connection to teammates. It addresses the feeling that others are truly co-located and responsive within the virtual space.
@@
Several articles emphasized that social presence is a foundational social-cognitive state that enables key collaborative behaviors such as turn-taking, mutual coordination, and shared understanding (Olaosebikan et al., 2022; Moharana et al., 2022; Dey et al., 2024; Kimmel et al., 2023; Wei et al., 2022; Weissker et al., 2023; Chastine et al., 2005). A key component of social presence in VR is mutual attention, which is the ability to perceive others’ attention and actions in real-time (Darbar et al., 2024; Bovo et al., 2022; Ide et al., 2020; Weissker et al., 2023). For example, Darbar et al. (Darbar et al., 2024) introduced ‘GazeMolVR,’ a system for molecular visualization in collaborative VR that incorporated shared eye-gaze cues. Their findings showed that real-time gaze sharing improved the feeling of connectedness and eased conversational flow. When collaborators saw others’ gaze, they adjusted their own attention more fluidly, making the interaction feel more natural and co-present (Darbar et al., 2024).
@@
The articles also demonstrate that social presence is strongly connected to nonverbal awareness and emotional connection. When avatars or systems convey cues such as gaze direction, pointing, or simple gestures, people can more easily interpret each other’s intentions (Bovo et al., 2022; Ide et al., 2020). For example, studies such as Bovo et al. (Bovo et al., 2022) and Ide et al. (Ide et al., 2020) implemented visual aids, including cones of vision and symbolic gestures, to help dyads perceive their collaborators’ focus and actions. These features not only added expressiveness but also enhanced social presence by increasing mutual visibility and acknowledgment, helping users feel noticed and understood by others.
@@
One article highlighted that social presence enhances coordination during multi-party interactions (Adkins et al., 2024). In their evaluation of audio and interaction modalities, Adkins et al. (Adkins et al., 2024) found that dyads using spatial audio, compared to non-spatial conditions, experienced smoother turn-taking and reported feeling more “in sync” with one another. These findings suggest that social presence is shaped not only by visual or avatar-based cues but also by auditory and behavioral alignment that supports dyadic coordination.
@@
One last article noted that weak or underdeveloped social presence negatively affects collaboration. In their evaluation of a co-design task, Mei et al. (Mei et al., 2021) found that the absence of avatar reactivity and expressive behavior made interactions feel more like a phone call rather than a shared, embodied collaboration. This diminished sense of social presence reduced user engagement and made communication feel less natural, highlighting the importance of social presence in enabling teamwork.
@@
In summary, these articles demonstrate that social presence serves as the social glue that makes collaboration in Social VR feel natural and connected. Ultimately, social presence enables members to perceive and respond to one another’s actions in real time, supporting cohesive and authentically shared collaboration in VR.
@@
This sub-theme includes articles that investigate how users evaluate the authenticity, seriousness, and naturalness of their collaborative experiences in Social VR. The focus is on whether immersive VR environments support interactions that are not only technically possible but also feel socially valid, natural, and meaningful. Beyond the technical functionality, these articles highlight that users carry expectations shaped by physical-world norms, emotional cues, and social dynamics, which influence how “real” collaboration feels in Social VR (Torro and Pirkkalainen, 2023; Maloney and Freeman, 2020; Osborne et al., 2023).
@@
Building on this, Maloney et al. (Maloney and Freeman, 2020) interviewed long-term Social VR users to understand what made virtual interactions feel meaningful. Participants reported that their collaborations felt more authentic when gestures, feedback, and shared activities aligned with everyday social norms. Many of them expressed frustration with unrealistic avatar features—such as exaggerated cartoon-like hands or faces that could not convey realistic emotions—which made the interactions feel artificial and less socially authentic. What users desired was not just visual realism, but interaction fidelity, which translated into responsiveness and expressiveness that matched the tone and rhythm of real-world collaboration.
Other articles in this sub-theme indicated that misaligned avatar behavior can disrupt the authenticity of collaboration. For example, Osborne et al. (Osborne et al., 2023) noted that technical glitches, such as avatar duplication or unnatural expressions, made interactions feel artificial and awkward. In contrast, when avatars moved and responded in socially expected ways, participants perceived the collaboration as smooth and believable, even when the avatars were abstract in design (Osborne et al., 2023).
@@
One study examined how to reinforce the role of realism in a Social VR workplace by eliminating fantastical elements. Park et al. (Park et al., 2024) removed the teleportation feature and required users to navigate through halls and elevators, mirroring how people would walk in a physical office. While this approach increased social encounters and enhanced users’ social presence, it also introduced usability challenges, revealing a tension between realism and practicality (Park et al., 2024). The findings of this article echo broader concerns about the necessary levels of realism to foster authentic collaboration without compromising the unique affordances of VR.
Lastly, one article (Dey et al., 2024) demonstrated that Social VR participants who first engaged in an emotionally intense solo experience approached a subsequent group task with more seriousness and coordination. The realism of the initial simulation made the following collaboration feel more consequential and authentic. Dey et al.’s findings (Dey et al., 2024) suggest that initial VR experiences can reshape how users’ perceptions of group work in Social VR, making virtual collaboration feel more meaningful, important, and socially grounded.
In summary, these articles demonstrate that when Social VR users feel that their actions have an impact and that their collaborators are genuinely engaged, the experience transcends the simulation and becomes a psychologically meaningful, socially credible form of collaboration.
@@
This sub-theme examines how trust, an essential yet often overlooked foundation of effective collaboration, is developed in Social VR. While trust in co-located settings often emerges through informal interactions and subtle nonverbal cues, trust-building mechanisms are often limited or absent in remote communication. The reviewed articles show that Social VR can restore some of these cues through embodied interaction, avatar behavior, and shared spatial context (Torro and Pirkkalainen, 2023; Yassien et al., 2020).
Torro et al. (Torro and Pirkkalainen, 2023) proposed a design theory explaining how Social VR can support trust-building in virtual teams. Grounded in social exchange theory—which emphasizes reciprocal interactions as the foundation of relationships (Cook et al., 2013)—they proposed that trust emerges through two key pathways in Social VR: a peripheral route, driven by first impressions (e.g., avatar appearance, social cues), and a central route, based on deeper assessments of a partner’s competence, integrity, and goodwill (Torro and Pirkkalainen, 2023). For example, avatars that mimic natural gestures, proximity-based conversations, and realistic movement can help promote social bonding. Moreover, features like virtual break rooms with spatial properties can support casual, unstructured interactions, which are often absent in structured video calls or text-based platforms.
Yassien et al. (Yassien et al., 2020) highlighted another key dimension of trust in Social VR: personal space. This article found that users felt safer and more connected when their virtual personal space was respected. Trust was undermined when avatars came too close or moved in ways that felt erratic or unpredictable, causing discomfort and breaking the sense of social ease. The researchers recommended designing Social VR applications with mechanisms to prevent collisions and unwanted proximity between avatars, reinforcing the importance of comfort and perceived safety.
Together, these two articles suggest that trust in Social VR is supported through both social signaling and environmental design, including expressive avatars, respectful personal space, and opportunities for informal interaction.
@@
This sub-theme focuses on social engagement, the extent to which collaborators in Social VR remain emotionally and communicatively connected during a shared task (Xu et al., 2023; Yang et al., 2022b). Social engagement goes beyond mere attention or performance; it reflects how present, expressive, and emotionally involved individuals feel in relation to their teammates (Xu et al., 2023; Yang et al., 2022b). The reviewed articles identified design strategies that support behavioral cues of engagement and help sustain interpersonal connections throughout the Social VR collaboration (Le Tarnec et al., 2024; Xu et al., 2023; Moore et al., 2019; Moharana et al., 2023; Gumilar et al., 2022; Duchowski et al., 2004).
One strategy identified in this sub-theme involves enhancing avatar behaviors to foster social engagement. Le et al. (Le Tarnec et al., 2024) introduced an augmented avatar system that generated subtle nonverbal cues—such as head nods and posture shifts—even when users did not perform them directly. These enhancements made collaborators appear more responsive and involved (Le Tarnec et al., 2024). Similarly, Xu et al. (Xu et al., 2023) studied older adults co-planning virtual travel experiences using VR and found that social engagement deepened when participants shared personal memories and life stories related to the destinations. These moments of self-disclosure fostered warmth and mutual interest, highlighting that social engagement can be supported not only through interface design but also through content that invites personal connection and emotional expression.
Another article explored physiological synchrony as a deeper measure of engagement. Moharana et al. (Moharana et al., 2023) tracked heart rate patterns during a collaborative memory game and found that dyads with higher physiological synchrony reported greater feelings of connection, presence, and workflow. Their findings suggest that social engagement may operate at both cognitive and emotional levels and that VR environments can help capture or even facilitate these more subtle social dynamics.
Taken together, these articles show that collaboration in Social VR extends beyond visual realism or task efficiency. It hinges on participants feeling seen, acknowledged, and emotionally connected, which are qualities that can be nurtured through expressive avatars, thoughtful interaction design, and moments that invite genuine interpersonal connections.